a) The sales representative here tells you they also have two floor plans available, but they only have 50 homes available. The representative tells you that floor plan #1 sells for $150,000 and floor plan #2 sells for $100,000. She also mentions that all the available houses combined are worth $7 million. Write a system of equations that illustrates this situation. Use x to represent floor plan #1 and y to represent floor plan #2




b) Use elimination method to determine how many houses with each floor plan are available. Explain how you arrived at your answer.

Answer by mananth(16946)

The total number of houses is 50
x+y=50..............1
The value of both plans = 7,000,000
150,000x+100,000y=7,000,000...........2
Plan I x Numbers
Plan II y Numbers

1 x + 1 y = 50 .............1
150000 x + 100000 y = 7000000 .............2
Eliminate y
multiply (1)by -100000
Multiply (2) by 1
-100000 x -100000 y = -5000000
150000 x + 100000 y = 7000000
Add the two equations
50000 x= 2000000
/ 50000
x = 40
plug value of x in (1)
1 x + 1 y = 50
40 + 1 y = 50
1 y = 50 -40
1 y = 10
y = 10
Plan I 40 Numbers
Plan II 10 Numbers

m.ananth@hotmail.ca